Homework Question: Daniel has a box shaped like a rectangular prism. It has a volume of 280 cubic inches. It is 8 inches long and 5 inches wide. Find the height of this box. Then find the Surface Area.
Answer by ankor@dixie-net.com(22740)

Daniel has a box shaped like a rectangular prism. It has a volume of 280 cubic inches.
It is 8 inches long and 5 inches wide.
Find the height of this box.
Using the volume equation L * W * H = vol
8 * 5 * h = 280
40h = 280
h = 280/40
h = 7
:
Then find the Surface Area.
S.A. = 2(L*W) + 2(LH) + 2(WH)
S.A. = 2(8*5) + 2(8*7) * 2(5*7)
I'll let you do the math here.